OpenStudy (kokoro):

(-3,-2),(5,3) slope intercept please help!

OpenStudy (anonymous):

Slope is y2 - y1 / x2 - x1 = 3 + 2 / 5 + 3 = 5/8

OpenStudy (anonymous):

slope intercept equation is given by y- y1 = m (x - x1) y + 2 = 5/8 (x + 3)

OpenStudy (kokoro):

Thank you but heres what confuses me

OpenStudy (kokoro):

When I convert it I have to distribute 5/8 for 3 but is my answer supposed to have decimals?

OpenStudy (anonymous):

o.625 replace this value with 5/8

OpenStudy (anonymous):

I find it easier to use Slope intercept form instead of point slope And 5/8 stays in fraction form

OpenStudy (kokoro):

Ah! Thank you Ganbil and Highhump

OpenStudy (kokoro):

Highjump* xD

OpenStudy (anonymous):

No problem :) OMG LOL. Nice typo

OpenStudy (anonymous):

:)